# EPW varieties as moduli spaces on ordinary GM surfaces and special GM threefolds

### v2

@@ 5.1. Birational model of the double dual EPW sextics
To achieve the isomorphism, one needs the description of the associated double dual EPW sextic in [18, Section 7.5].
At first, one notices that each line $L\subset S$ gives a Lagrangian submanifold $\mathbb{P}^{2}\cong(\mathbb{P}^{1})^{[2]}$ contained in $S^{[2]}$. In this case, one has a Mukai flop
@@
with respect to all these Lagrangian planes. The variety $\bar{X}_{S}$ is a subvariety of $\operatorname{Gr}(2,W)$ and the contraction $S^{[2]}\rightarrow\bar{X}_{S}$ is given by sending a point in $S^{[2]}$ to its spanned line in $\operatorname{Gr}(2,W)$.
@@
Also, one notices that a conic (including the degenerated ones) and a line in $S$ cannot meet exactly at a length $2$ subscheme due to Theorem 2.4. So the dual plane $\Sigma_{C}^{\vee}$ parameterizing lines in the plane $\Sigma_{C}$ spanned by $C$ can be realized as a subvariety in $X_{S}$.
@@
$\bm{p}_{S}\in Y^{\geq 3}_{A(S)^{\perp}}=\{[v]\in\mathbb{P}(V_{6}^{\vee})\,|\,\dim(A^{\perp}_{S}\cap\left(v\land\mathsf{\Lambda}^{2}V_{6}^{\vee}\right))\geq 3\}$
determined by the subspace $V_{5}\subset V_{6}$ according to [13, Theorem 3.10], and $Y^{\geq 3}_{A(S)^{\perp}}-\{\bm{p}_{S}\}$ is equal to the set of conics in $S$. The point $\bm{p}_{C}\in Y^{\geq 3}_{A(S)^{\perp}}$ corresponding to a conic $C\subset S$ is given by the unique 5-dimensional subspace $U_{5}\subset V_{6}$ determined by the linear span of $C$ in $\mathbb{P}(W)$.
Theorem 5.1 ([18, Theorem 7.7 and Remark 7.9]).
@@
The Mukai flop $g$ coincides with the flop described in [39, (4.2.4)]. Once $S$ does not contain lines and conics, the resolution $f^{+}$ coincides with the one in [39, (4.2.2)].
@@ 5.2. Exclusion of the non-flopping walls
Suppose otherwise that there such a vector $w=(r,\Delta,s)$, then one has has $(t+9)\Delta\cdot H=-40r$ and $\Delta^{2}=2r^{2}-2$. It follows
@@
and, after plug $s$ and $r$, it becomes
@@
which follows $\Delta\cdot H<0$. Moreover, one has
@@ 5.3. The wall-crossing of spherical factors
implying that $\Delta\cdot H+4r$ can not be an integer, which is absurd ∎
@@
Since (1,L,0)(1,-L,0) is a spherical vector, the stable HN factor $A$ of $\mathscr{O}_{S}(-L)$ is also spherical due to Proposition 3.19 (or directly [11, Lemma 12.2]). So the Hodge index theorem gives
@@ 5.4. The flopping walls and identification
According to Theorem 3.18, a flopping wall for it can only be induced by a pairs of vectors $w$ and $w^{\prime}$ in $\textbf{{H}}_{W}$ such that $\langle w,w\rangle=\langle w^{\prime},w^{\prime}\rangle=-2,\langle v_{1},w\rangle=\langle v_{1},w^{\prime}\rangle=1$ and $w^{\prime}+w=v_{1}$. One writes $w=(r,\Delta,s)$, then $(t+9)\Delta\cdot H=-40r+20$ and $\Delta^{2}=2r^{2}-2r-2$. It follows
@@
The divisor class $\Delta$ in $(0,\Delta,-1)$ is in fact the class of a line in $S$ (see, for example, the discussions in [24, Section 2.1.4]) and the class $\Delta$ in $(1,\Delta,0)$ is L-L for some line $L$ in $S$.
@@
According to Theorem 3.18, a flopping wall for $\textsf{{M}}_{\sigma_{S}}(v_{1})$ is induced by a pairs of vectors $w$ and $w^{\prime}$ in $\textbf{{H}}_{W}$ such that $\langle w,w\rangle=\langle w^{\prime},w^{\prime}\rangle=-2,\langle v_{1},w\rangle=\langle v_{1},w^{\prime}\rangle=1$ and $w^{\prime}+w=v_{1}$. One writes $w=(r,\Delta,s)$, then $\Delta\cdot H=-4r+2$ and also $\Delta^{2}=2r^{2}-2r-2$. It follows
@@
Also, one can see that the class $\Delta$ in $w=(1,-\Delta,0)$ and $w^{\prime}=(0,\Delta,-1)$ with $\Delta\cdot H=2$ and $\Delta^{2}=-2$ is in fact the class of a conic $C$ in $S$ (see for example [24, Section 2.1.4]). The corresponding flopping wall is induced by the short exact sequences
@@
where $Z\subset S$ is a length $2$ subscheme spanning a line not in $S$. Indeed, such a short exact sequence follows from the fact that the ideal sheaf $\mathscr{I}_{L/M_{S}}$ of every line $L\subset M_{S}$ can be written as the cokernel of a unique map $\mathscr{U}_{G}|_{M_{S}}\rightarrow\mathscr{V}_{G}|_{M_{S}}$, and each such map has the ideal sheaf of a line as the cokernel. One just takes the restriction onto $S\subset M_{S}$. The line spanned by $Z$ is not in $S$ because such a length two subscheme $Z$ has been already flopped in Proposition 5.13. So one contracts the Fano variety $F(M_{S})$ of lines, and therefore recovers $f^{+}\colon X_{S}\rightarrow\tilde{Y}_{A(S)^{\perp}}^{\geq 1}$. ∎
@@
Moreover, the involution on $\textsf{{M}}_{\sigma_{S}}(v_{1})$ induced by the group action $\mathfrak{S}^{\vee}_{2}$ coincides with the covering involution on the double dual EPW sextics as the Hodge isometries of the corresponding birational involutions on $X_{S}$ are the same (see [33, Section 6.5] for the generic case). A point-wise description of the involution can be found in [18, Remark 7.8]. Also, the fixed locus $\textsf{{M}}_{\sigma_{S}}(v_{1})_{\mathfrak{S}^{\vee}_{2}}$ of the involution is exactly the ramification locus $Y^{\geq 2}_{A(S)^{\perp}}$.
@@ 6.1. Double EPW sextics as moduli spaces
It is shown in [33, Section 6.6] that the moduli space $\textsf{{M}}_{\sigma_{S}}(v_{2})$ is isomorphic to the double EPW sextics once $S$ does not contain any lines. Here one can generalize it using Theorem 1.1 and knowledge on the moduli space and the period of double EPW sextics. To continue, let us recall some notations from [40].
The moduli space of double EPW sextics is defined as the GIT quotient
where $\operatorname{LG}(\mathsf{\Lambda}^{3}V_{6})\subset\operatorname{Gr}(10,\land^{3}V)$ is the symplectic Grassmannian parameterizing subspaces which are Lagrangian for the symplectic form given by wedge-product. The period domain $\mathbb{D}_{\Lambda}^{\textup{BB}}$ is defined by the Baily–Borel compactification of the quotient $\mathbb{D}_{\Lambda}$ of a bounded symmetric domain of Type IV by the action of an arithmetic group (see [40, Section 4.2] for details).
Then there exists a rational map
$\mathfrak{p}\colon\mathfrak{M}\dashrightarrow\mathbb{D}_{\Lambda}^{\textup{BB}}$
defined by the hyperkähler period on the open locus where double EPW sextic is a hyperkähler fourfold. According to [40], the rational map is a regular at the open locus $\mathfrak{M}$-$\mathfrak{R}$, where $\mathfrak{R}$ is determined by the Lagrangian subspaces containing decomposable elements.
The mapping $A\mapsto A^{\perp}$ provide an involution $\tau$ on the moduli space $\mathfrak{M}$-$\mathfrak{R}$ which is compatible with the period map $\mathfrak{p}$ i.e. $\mathfrak{p}\circ\tau=\bar{\tau}\circ\mathfrak{p}$ for some involution $\bar{\tau}$ on $\mathbb{D}_{\Lambda}^{\textup{BB}}$.
Proposition 6.1 ([40, Section 4.3 and Theorem 5.1]).
The ramification locus of the involution $\bar{\tau}$ does not intersect with the image $\mathfrak{p}(\mathfrak{M}-\mathfrak{R})$ in $\mathbb{D}_{\Lambda}^{\textup{BB}}$.
@@
The moduli space $\textsf{{M}}_{\sigma_{S}}(v_{2})$ is isomorphic to the double EPW sextics.
@@
The generic case has been verified in [33, Section 6.6], so it remains to prove the statement when $Y^{\geq 3}_{A(S)}\neq 0$. In this case, the special Gushel–Mukai threefold $X$ over $S$ admits a period dual $X^{\prime}$ which is also a special Gushel–Mukai threefold by [31, Lemma 3.8]. Then it follows from [32] that there exists an equivalence ${\textbf{{Ku}}}(X)\cong{\textbf{{Ku}}}(X^{\prime})$. This equivalence descends to an equivalence ${\textbf{{D}}}^{b}(S)\cong{\textbf{{D}}}^{b}(S^{\prime})$ for the strongly smooth Gushel–Mukai surface $S$ associated with the threefold $X^{\prime}$ which is equivariant with the $\mathfrak{S}_{2}^{\vee}$-group actions on both side. So the stability condition $\sigma_{S}$ is sent to $\sigma_{S^{\prime}}$ and the invariant lattice $\mathbb{Z}[v_{1}]\oplus\mathbb{Z}[v_{2}]$ is sent to $\mathbb{Z}[v^{\prime}_{1}]\oplus\mathbb{Z}[v^{\prime}_{2}]$.
In particular, the equivalence either sends $v_{1}$ to $v^{\prime}_{1}$ and $v_{2}$ to $v^{\prime}_{2}$ up to a sign or exchanges them. The first case will lead to $\tilde{Y}^{\geq 1}_{A(S)^{\perp}}\cong\textsf{{M}}_{\sigma_{S}}(v_{1})\cong\textsf{{M}}_{\sigma_{S^{\prime}}}(v^{\prime}_{1})\cong\tilde{Y}^{\geq 1}_{A(S)}$ and then $A(S)\cong A(S)^{\perp}$ due to Theorem 2.10 which will violate the previous proposition. So one concludes. ∎
@@
Consider two special Gushel–Mukai threefolds $X$ and $X^{\prime}$. Then $X$ and $X^{\prime}$ are period partner if and only if one has an equivalence ${\textbf{{Ku}}}(X)\cong{\textbf{{Ku}}}(X^{\prime})$ sending $\kappa_{1}$ to $\kappa^{\prime}_{1}$ and $\kappa_{2}$ to $\kappa^{\prime}_{2}$ up to a sign; $X$ and $X^{\prime}$ are period dual if and only if one has an equivalence ${\textbf{{Ku}}}(X)\cong{\textbf{{Ku}}}(X^{\prime})$ sending $\kappa_{1}$ to $\kappa^{\prime}_{2}$ and $\kappa_{1}$ to $\kappa^{\prime}_{2}$ up to a sign.
@@ 6.2. Double dual EPW surfaces as moduli spaces
The statements in this subsection are actually just corollaries of [18, Theorem 7.3] and [28, Theorem 7.12]. The first one describes the associated double dual EPW surfaces, and second one describes the moduli spaces.
Here we only elaborate on the case for a special Gushel–Mukai threefold $X$ over $S$ and the starting point is the Hilbert scheme $C(X)$ of conics on $X$. The surface $C(X)$ has two components $C_{\sigma}(X)$ and $C_{0}(X)$ such that $C_{\sigma}(X)\cong F(M_{S})\cong\mathbb{P}^{2}$ parameterizes the preimage of lines on $M_{S}$. The other component $C_{0}(X)$ is isomorphic to the blow up of the double dual EPW surface associated with $X$ at the distinguished point $\tilde{\bm{p}}_{S}$ according to [18, Theorem 7.3]. These components intersect transversely along the exceptional curve on $C_{0}(X)$.
Due to [28, Theorem 7.12], the moduli space $\textsf{{M}}_{\sigma_{X}}(-\kappa_{1})$ is a contraction from $C(X)$ such that the component $C_{\sigma}(X)$ is contracted to the distinguished singular point $\tilde{\bm{p}}_{S}$. It follows that the space $\textsf{{M}}_{\sigma_{X}}(-\kappa_{1})$ is nothing but the double dual EPW surface associated with $X$.
@@
The isomorphism $\textsf{{M}}_{\sigma_{X}}(-\kappa_{1})\cong\tilde{Y}^{\geq 2}_{A(S)^{\perp}}$ follows from [18, 28] as above. The identification of these branched double covers is similar to [33, Proposition 5.20]. One can do this because the involution on $C(X)$ induced by the branched cover $X\rightarrow M_{S}$ coincides with the one induced by the natural covering involution on $\tilde{Y}^{\geq 2}_{A(S)^{\perp}}$ according to [18, Section 7.3]. In particular, the branched double cover for the double dual EPW surface $\tilde{Y}^{\geq 2}_{A(S)^{\perp}}$ can be described in terms of conics on an open subset and therefore be identified. ∎
@@
It is again a direct application of [18, Theorem 7.3] and [28, Theorem 7.12]. ∎

### v3

@@ 5.1. Birational model of the double dual EPW sextics
To achieve the isomorphism, one needs the description of the associated double dual EPW sextic in [21, Section 7.5].
At first, one notices that each line $L\subset S$ gives a Lagrangian submanifold $\mathbb{P}^{2}\cong(\mathbb{P}^{1})^{[2]}$ in the hyperkähler manifold $S^{[2]}$. Then, one has a Mukai flop
@@
of these Lagrangian planes in $S^{[2]}$. The variety $\bar{X}_{S}$ is a subvariety of $\operatorname{Gr}(2,W)$ and the contraction $S^{[2]}\rightarrow\bar{X}_{S}$ is given by sending a point in $S^{[2]}$ to its spanned line in $\operatorname{Gr}(2,W)$.
@@
Also, one notices that a conic (including the degenerate ones) and a line in $S$ cannot meet exactly at a length $2$ subscheme due to Theorem 2.4. So the dual plane $\Sigma_{C}^{\vee}$ parameterizing lines in the plane $\Sigma_{C}$ spanned by $C$ can be realized as a subvariety in $X_{S}$.
@@
$\bm{p}_{S}\in Y^{\geq 3}_{A(S)^{\perp}}=\{[v]\in\mathbb{P}(V_{6}^{\vee})\,|\,\dim(A(S)^{\perp}\cap\left(v\land\mathsf{\Lambda}^{2}V_{6}^{\vee}\right))\geq 3\}$
determined by the subspace $V_{5}\subset V_{6}$ according to [16, Theorem 3.10], and $Y^{\geq 3}_{A(S)^{\perp}}-\{\bm{p}_{S}\}$ is equal to the set of conics in $S$. The point $\bm{p}_{C}\in Y^{\geq 3}_{A(S)^{\perp}}$ corresponding to a conic $C\subset S$ is given by the unique 5-dimensional subspace $U_{5}\subset V_{6}$ determined by the linear span of $C$ in $\mathbb{P}(W)$.
Theorem 5.1 ([21, Theorem 7.7 and Remark 7.9]).
@@
The Mukai flop $g$ coincides with the flop described in [43, (4.2.4)]. Once $S$ does not contain lines and conics, the resolution $f^{+}$ coincides with the one in [43, (4.2.2)].
@@ 5.2. Exclusion of the non-flopping walls
Suppose otherwise that there is such a vector $w=(r,\Delta,s)$, then one has $(t+9)\Delta\cdot H=-40r$ and $\Delta^{2}=2r^{2}-2$. It follows
@@
and, after plugging $s$ and $r$, it becomes
@@
which implies $\Delta\cdot H<0$. Moreover, one has
@@ 5.3. The wall-crossing of spherical factors
implying that $\Delta\cdot H+4r$ can not be an integer, which is absurd. ∎
@@
Since (1,L,0)(1,-L,0) is a spherical vector, the stable HN factor $A$ of $\mathscr{O}_{S}(-L)$ is also spherical due to Proposition 3.19 (or directly [14, Lemma 12.2]). So the Hodge index theorem gives
@@ 5.4. The flopping walls and identification
According to Theorem 3.18, a flopping wall for it can only be induced by a pair of vectors $w$ and $w^{\prime}$ in $\textbf{{H}}_{W}$ such that $\langle w,w\rangle=\langle w^{\prime},w^{\prime}\rangle=-2,\langle v_{1},w\rangle=\langle v_{1},w^{\prime}\rangle=1$ and $w^{\prime}+w=v_{1}$. One writes $w=(r,\Delta,s)$, then $(t+9)\Delta\cdot H=-40r+20$ and $\Delta^{2}=2r^{2}-2r-2$. It follows
@@
The divisor class $\Delta$ in $(0,\Delta,-1)$ is in fact the class of a line in $S$ (see, for example, the discussions in [27, Section 2.1.4]) and the class $\Delta$ in $(1,\Delta,0)$ is L-L for some line $L$ in $S$.
@@
According to Theorem 3.18, a flopping wall for $\textsf{{M}}_{\sigma_{S}}(v_{1})$ is induced by a pair of vectors $w$ and $w^{\prime}$ in $\textbf{{H}}_{W}$ such that $\langle w,w\rangle=\langle w^{\prime},w^{\prime}\rangle=-2,\langle v_{1},w\rangle=\langle v_{1},w^{\prime}\rangle=1$ and $w^{\prime}+w=v_{1}$. One writes $w=(r,\Delta,s)$, then $\Delta\cdot H=-4r+2$ and also $\Delta^{2}=2r^{2}-2r-2$. It follows
@@
Also, one can see that the class $\Delta$ in $w=(1,-\Delta,0)$ and $w^{\prime}=(0,\Delta,-1)$ with $\Delta\cdot H=2$ and $\Delta^{2}=-2$ is in fact the class of a conic $C$ in $S$ (see for example [27, Section 2.1.4]). The corresponding flopping wall is induced by the short exact sequences
@@
where $Z\subset S$ is a length $2$ subscheme spanning a line not in $S$. Indeed, such a short exact sequence follows from the fact that the ideal sheaf $\mathscr{I}_{L/M_{S}}$ of every line $L\subset M_{S}$ can be written as the cokernel of a unique map $\mathscr{U}_{G}|_{M_{S}}\rightarrow\mathscr{V}_{G}|_{M_{S}}$, and each such map has the ideal sheaf of a line as the cokernel. One just takes the restriction onto $S\subset M_{S}$. The line spanned by $Z$ is not in $S$ because such a length two subscheme $Z$ has already been flopped in Proposition 5.13. So one contracts the Fano variety $F(M_{S})$ of lines, and therefore recovers $f^{+}\colon X_{S}\rightarrow\tilde{Y}_{A(S)^{\perp}}^{\geq 1}$. ∎
@@
Moreover, the involution on $\textsf{{M}}_{\sigma_{S}}(v_{1})$ induced by the group action $\mathfrak{S}^{\vee}_{2}$ coincides with the covering involution on the double dual EPW sextic as the Hodge isometries of the corresponding birational involutions on $X_{S}$ are the same (see [37, Section 6.5] for the generic case). A point-wise description of the involution can be found in [21, Remark 7.8]. Also, the fixed locus $\textsf{{M}}_{\sigma_{S}}(v_{1})_{\mathfrak{S}^{\vee}_{2}}$ of the involution is exactly the ramification locus $Y^{\geq 2}_{A(S)^{\perp}}$.
@@ 6.1. Double EPW sextics as moduli spaces
It is shown in [37, Section 6.6] that the moduli space $\textsf{{M}}_{\sigma_{S}}(v_{2})$ is isomorphic to the double EPW sextic once $S$ has Picard rank one. Here let us extend the result to the case that $Y_{A(S)}^{\geq 3}=\varnothing$. According to [11, Theorem 3.3], it means that the surface $S$ contains no lines or quintic elliptic curves.
Proposition 6.1.
Suppose that $(S,H)$ is a strongly smooth Gushel–Mukai surface containing no lines or quintic elliptic curves, then one has $\textsf{{M}}_{\sigma_{S}}(v_{2})\cong M_{H}(-v_{2})\cong\tilde{Y}_{A(S)}^{\geq 1}$.
Proof.
It suffices to show that $\sigma_{t}$ is $v_{2}$-generic for any $t\geq 1$. Otherwise, one chooses the largest $t\geq 1$ such that $\sigma_{t}$ is on a wall with respect to $v_{2}$. Then a strictly $\sigma_{t}$-semistable object in heart with Mukai vector $v_{2}$ has the form $\mathscr{G}[1]$ with $\mathscr{G}\in\textsf{{M}}_{H}(-v_{2})=M_{H}(2,-H,2)$.
One finds a short exact sequence $0\rightarrow A\rightarrow\mathscr{G}[1]\rightarrow B\rightarrow 0$ with the same phase with $A$ stable and $\mu(A)=\mu(B)$. At first, one can see $r\geq-1$ for $v(A)=(r,\Delta,s)$ from the long exact sequence of cohomology sheaves $0\rightarrow\mathscr{H}^{-1}(A)\rightarrow\mathscr{G}\rightarrow\mathscr{H}^{-1}(B)\rightarrow\mathscr{H}^{0}(A)\rightarrow 0$.
Also, one can obtain
$\Delta\cdot H=1-4r\quad\textup{and}\quad s=\frac{t+4}{5}r+\frac{t-1}{5}$
through $\mu(A)=\mu(B)=\mu(\mathscr{G}[1])<+\infty$. Then one can see
$(4r-1)^{2}=(\Delta\cdot H)^{2}\geq 10\Delta^{2}\geq 20rs-20=4(t+4)r^{2}+(t-1)r-20$
using the Hodge index theorem and Mukai lemma. It follows
$\frac{21}{5}\geq(tr+1)(r+1)$
and one can see $r\leq 1$ from this inequality and $t\geq 1$.
Suppose that $r=1$, then $\Delta\cdot H=-3$ and $9\geq 10\Delta^{2}\geq 5t-23\geq-18$. It follows $\Delta^{2}=0$, but then the class $\Delta$ will violate Proposition 2.4.
Suppose that $r=0$, then $\Delta\cdot H=1$ and $1\geq 10\Delta^{2}\geq-20$. It follows $\Delta^{2}=-2$, but then the class $\Delta$ is a line which is impossible.
Suppose that r=1r=-1, then $\Delta\cdot H=5$ and $25\geq 10\Delta^{2}\geq 0$. It follows $\Delta^{2}=2$, as $S$ also does not contain quintic elliptic curves. However, in this case the class $D:=H-2\Delta$ is a degree zero effective class, which is impossible as $H$ is ample. ∎
Consequently, one can obtain the following statement.
@@
The moduli space $\textsf{{M}}_{\sigma_{S}}(v_{2})$ is isomorphic to the double EPW sextic.
@@
The generic case has been verified in Proposition 6.1, so it remains to prove the statement when $Y^{\geq 3}_{A(S)}\neq\varnothing$. In this case, the special Gushel–Mukai threefold $X$ over $S$ admits a period dual $X^{\prime}$ which is also a special Gushel–Mukai threefold by [34, Lemma 3.8]. Then it follows from [35] that there exists an equivalence ${\textbf{{Ku}}}(X)\cong{\textbf{{Ku}}}(X^{\prime})$. This equivalence descends to an equivalence ${\textbf{{D}}}^{b}(S)\cong{\textbf{{D}}}^{b}(S^{\prime})$ for the strongly smooth Gushel–Mukai surface $S^{\prime}$ associated with the threefold $X^{\prime}$ which is equivariant with the $\mathfrak{S}_{2}^{\vee}$-group actions on both side. So the stability condition $\sigma_{S}$ is sent to $\sigma_{S^{\prime}}$ and the invariant lattice $\mathbb{Z}[v_{1}]\oplus\mathbb{Z}[v_{2}]$ is sent to $\mathbb{Z}[v^{\prime}_{1}]\oplus\mathbb{Z}[v^{\prime}_{2}]$.
In particular, the equivalence either sends $v_{1}$ to $v^{\prime}_{1}$ and $v_{2}$ to $v^{\prime}_{2}$ up to a sign or exchanges them by structure of the lattices. The first case will lead to $\tilde{Y}^{\geq 1}_{A(S)^{\perp}}\cong\textsf{{M}}_{\sigma_{S}}(v_{1})\cong\textsf{{M}}_{\sigma_{S^{\prime}}}(v^{\prime}_{1})\cong\tilde{Y}^{\geq 1}_{A(S)}$ which will violate Theorem 2.10 and Proposition A.3. So one concludes. ∎
@@
Consider two special Gushel–Mukai threefolds $X$ and $X^{\prime}$. Then $X$ and $X^{\prime}$ are period partner if and only if one has an equivalence ${\textbf{{Ku}}}(X)\cong{\textbf{{Ku}}}(X^{\prime})$ sending $\kappa_{1}$ to $\kappa^{\prime}_{1}$ and $\kappa_{2}$ to $\kappa^{\prime}_{2}$ up to a sign; $X$ and $X^{\prime}$ are period dual if and only if one has an equivalence ${\textbf{{Ku}}}(X)\cong{\textbf{{Ku}}}(X^{\prime})$ sending $\kappa_{1}$ to $\kappa^{\prime}_{2}$ and $\kappa_{2}$ to $\kappa^{\prime}_{1}$ up to a sign.
@@ 6.2. Double dual EPW surfaces as moduli spaces
The statements in this subsection are actually just corollaries of [21, Theorem 7.3] and [31, Theorem 7.12]. The first one describes the associated double dual EPW surfaces, and the second one describes the moduli spaces.
Here we only elaborate on the case for a special Gushel–Mukai threefold $X$ over $S$ and the starting point is the Hilbert scheme $C(X)$ of conics on $X$. The surface $C(X)$ has two components $C_{\sigma}(X)$ and $C_{0}(X)$ such that $C_{\sigma}(X)\cong F(M_{S})\cong\mathbb{P}^{2}$ parameterizes the preimage of lines on $M_{S}$. The other component $C_{0}(X)$ is isomorphic to the blow up of the double dual EPW surface associated with $X$ at the distinguished point $\tilde{\bm{p}}_{S}$ according to [21, Theorem 7.3]. These components intersect transversely along the exceptional curve on $C_{0}(X)$.
Due to [31, Theorem 7.12], the moduli space $\textsf{{M}}_{\sigma_{X}}(-\kappa_{1})$ is a contraction from $C(X)$ such that the component $C_{\sigma}(X)$ is contracted to the distinguished singular point $\tilde{\bm{p}}_{S}$. It follows that the space $\textsf{{M}}_{\sigma_{X}}(-\kappa_{1})$ is nothing but the double dual EPW surface associated with $X$.
@@
The isomorphism $\textsf{{M}}_{\sigma_{X}}(-\kappa_{1})\cong\tilde{Y}^{\geq 2}_{A(S)^{\perp}}$ follows from [21, 31] as above. The identification of these branched double covers is similar to [37, Proposition 5.20]. One can do this because the involution on $C(X)$ induced by the branched cover $X\rightarrow M_{S}$ coincides with the one induced by the natural covering involution on $\tilde{Y}^{\geq 2}_{A(S)^{\perp}}$ according to [21, Section 7.3]. In particular, the branched double cover for the double dual EPW surface $\tilde{Y}^{\geq 2}_{A(S)^{\perp}}$ can be described in terms of conics on an open subset and therefore be identified. ∎
@@
It is again a direct application of [21, Theorem 7.3] and [31, Theorem 7.12]. ∎